# Finite temperature detection of quantum critical points via internal quantum teleportation

### v2

@@ V Discussion
Finally, we should also contrast the present proposal with the standard way of studying QPTs from the experimental point of view. In the standard way, we must measure the one- and two-point correlation functions using, for instance, neutron scattering techniques. The non-analytic properties of these one- and two-point correlation functions at the QCP marks a QPT. In the present proposal, we need only to measure one-point correlation functions after “disturbing” the system in a specific way. Indeed, to reconstruct the state describing his qubit after the teleportation protocol (the “disturbance” in the spin chain), Bob only needs one-point correlation functions. Any single qubit density matrix is completely characterized once we know all the one-point correlation functions $\left\langle\sigma^{x}\right\rangle,\left\langle\sigma^{y}\right\rangle$, and $\left\langle\sigma^{z}\right\rangle$ nie00 . And in all the models investigated here, we only need one correlation function, $\left\langle\sigma^{z}\right\rangle$ to be specific. This is true since the qubit with Bob after the teleportation protocol is diagonal in the standard basis [see Eqs. (48)-(57)]. In other words, we are trading the measurement of two-point correlation functions to experimentally determine a QCP for the measurement of only one one-point correlation function after disturbing the system in a very particular way, namely, after applying all the gates associated with the internal quantum teleportation protocol.
@@ VI Conclusion
We simplified the teleportation based quantum critical point (QCP) detector of Ref. pav23 in a very important way, reducing the technical demands to its possible experimental realization. Instead of using an external qubit from the spin chain as the input state to be teleported to the chain, in this work we employed a qubit within the spin chain itself. Several spin-1/2 chain models in the thermodynamic limit were used to test the efficacy of this new approach in detecting a QCP at zero and non-zero temperatures. We showed that whenever the spin chain is immersed in an external magnetic field, the present “internal” teleportation based QCP detector retains all the remarkable characteristics of the “external” teleportation based QCP detector of Ref. pav23 . In particular, we showed that the present strategy detects the QCPs for the XXZ model in a longitudinal field, the Ising transverse model, the isotropic XX model in a transverse field, and the anisotropic XY model in a transverse field.
@@
For the several models investigated here, we observed that at $T=0$ the maximum mean fidelity $\overline{\mathcal{F}}$ and the minimum mean trace distance $\overline{\mathcal{D}}$ between the teleported and the input states possess a cusp or an inflection point exactly the QCPs. When we increase $T$, these cusps are smoothed out and both the cusps and the inflection points are displaced from the exact locations of the QCPs. At finite temperatures, these cusps and inflection points manifest themselves in a high value for the magnitudes of the first and second order derivatives of $\overline{\mathcal{F}}$ and $\overline{\mathcal{D}}$ around the exact locations of the QCPs. We also showed that for small values of temperature, the extrema of these derivatives lie close together and by taking the zero temperature limit we are able to correctly predict the $T=0$ location of the QCP.
@@
Furthermore, the internal teleportation based QCP detector share the same features of the external one and of the quantum discord, when the latter is also employed as a tool to detect QCPs wer10b . These tools can be used without the knowledge of the order parameter associated with the quantum phase transition being investigated and they are very resilient to the increase of temperature, giving us useful information at finite $T$ that allows us to predict the exact value of the QCP at $T=0$. On top of that, and contrary to the quantum discord, the QCP detectors here developed and in Ref. pav23 have a straightforward experimental meaning and are scalable to high spin systems pav23 .
@@
We want to end this work calling attention to a feature already highlighted in Ref. pav23 and that is also observed here. This feature is related to the fact that for each model investigated in this work we have a distinctive functional form for $\overline{\mathcal{F}}$ and $\overline{\mathcal{D}}$. For some models we have extra cusps for these functions that are not related to a quantum phase transition, for other models the QCP is detected via an inflection point of $\overline{\mathcal{F}}$ and $\overline{\mathcal{D}}$, and for others we have the QCP being detected by discontinuities in their derivatives. Putting it differently, the “fingerprint” of a quantum phase, a quantum phase transition, and the underlying model dictating these phases and transitions are unique. The functional forms of $\overline{\mathcal{F}}$ and $\overline{\mathcal{D}}$ as we drive the system around its phase space are specific for a given model. A systematic study $\overline{\mathcal{F}}$ and $\overline{\mathcal{D}}$ enables the proper detection of the QCP using finite $T$ data and, interestingly, also the discovery of the underlying spin chain model giving rise to those quantum phases and quantum phase transitions.

### v3

@@ V Discussion
We should also contrast the present proposal with the standard way of studying QPTs from the experimental point of view. In the standard way, we must measure the one- and two-point correlation functions using, for instance, neutron scattering techniques. The non-analytic properties of these one- and two-point correlation functions at the QCP mark a QPT. In the present proposal, we need only to measure one-point correlation functions after “disturbing” the system in a specific way. Indeed, to reconstruct the state describing his qubit after the teleportation protocol (the “disturbance” in the spin chain), Bob only needs one-point correlation functions. Any single qubit density matrix is completely characterized once we know all the one-point correlation functions $\left\langle\sigma^{x}\right\rangle,\left\langle\sigma^{y}\right\rangle$, and $\left\langle\sigma^{z}\right\rangle$ nie00 . And in all the models investigated here, we only need one correlation function, $\left\langle\sigma^{z}\right\rangle$ to be specific. This is true since the qubit with Bob after the teleportation protocol is diagonal in the standard basis [see Eqs. (48)-(57)]. In other words, we are trading the measurement of two-point correlation functions to experimentally determine a QCP for the measurement of only one one-point correlation function after disturbing the system in a very particular way, namely, after applying all the gates associated with the internal quantum teleportation protocol.
Finally, we want to emphasize that the previous paragraph in no way implies that the present proposal is experimentally simpler or easier than the standard approach to detect QCP, namely, the direct measurement of one- and two-point correlation functions. Actually, the internal teleportation based QCP detector can be applied to detect a QCP whether or not it is experimentally implemented as described in the previous paragraphs. The same standard experimental techniques to measure one- and two-point correlation functions serve as well. Once we have those correlation functions, we automatically obtain Eq. (20), from which the maximal mean fidelity $\overline{\mathcal{F}}$ and the minimum mean trace distance $\overline{\mathcal{D}}$ associated with the internal teleportation based QCP detector can be computed. And if we look at Eqs. (42) and (45), we realize that it is the non-linear dependence of $\overline{\mathcal{F}}$ and $\overline{\mathcal{D}}$ on the one- and two-point correlation functions that sets them apart from the external teleportation QCP detector, where the fidelity is a simple linear function of a specific two-point correlation function. We believe that this unique functional form of $\overline{\mathcal{F}}$ and $\overline{\mathcal{D}}$ is the reason of why the internal teleportation QCP detector outperforms the external one most of the time.
@@ VI Conclusion
We simplified the teleportation based quantum critical point (QCP) detector of Ref. pav23 in a very important way, reducing the technical demands to its possible experimental realization. Instead of using an external qubit from the spin chain as the input state to be teleported to the chain, in this work we employed a qubit within the spin chain itself. Several spin-1/2 chain models in the thermodynamic limit were used to test the efficacy of this new approach in detecting a QCP at zero and non-zero temperatures. We showed that whenever the spin chain is immersed in an external magnetic field, the present “internal” teleportation based QCP detector retains all the remarkable characteristics of the “external” teleportation based QCP detector of Ref. pav23 . In particular, we showed that the present strategy detects the QCPs for the XXZ model in a longitudinal field, the Ising transverse model, the isotropic XX model in a transverse field, and the anisotropic XY model in a transverse field. Note that we worked with these models because their QCPs were previously characterized by other methods. In this way we could properly compare the predicted location for the QCPs coming from the present tool with the well-known and established locations for those QCPs.
@@
For the several models investigated here, we observed that at $T=0$ the maximum mean fidelity $\overline{\mathcal{F}}$ and the minimum mean trace distance $\overline{\mathcal{D}}$ between the teleported and the input states possess a cusp or an inflection point exactly at the QCPs. When we increase $T$, these cusps are smoothed out and both the cusps and the inflection points are displaced from the exact locations of the QCPs. At finite temperatures, these cusps and inflection points manifest themselves in a high value for the magnitudes of the first and second order derivatives of $\overline{\mathcal{F}}$ and $\overline{\mathcal{D}}$ around the exact locations of the QCPs. We also showed that for small values of temperature, the extrema of these derivatives lie close together and by taking the zero temperature limit we are able to correctly predict the $T=0$ location of the QCP.
@@
Furthermore, the internal teleportation based QCP detector shares the same features of the external one and of the quantum discord, when the latter is also employed as a tool to detect QCPs wer10b . These tools can be used without the knowledge of the order parameter associated with the quantum phase transition being investigated and they are very resilient to the increase of temperature, giving us useful information at finite $T$ that allows us to predict the exact value of the QCP at $T=0$. On top of that, and contrary to the quantum discord, the QCP detectors here developed and in Ref. pav23 have a straightforward experimental meaning and are scalable to high spin systems pav23 .
@@
We want to end this work by calling attention to a feature already highlighted in Ref. pav23 and that is also observed here. This feature is related to the fact that for each model investigated in this work we have a distinctive functional form for $\overline{\mathcal{F}}$ and $\overline{\mathcal{D}}$. For some models we have extra cusps for these functions that are not related to a quantum phase transition, for other models the QCP is detected via an inflection point of $\overline{\mathcal{F}}$ and $\overline{\mathcal{D}}$, and for others we have the QCP being detected by discontinuities in their derivatives. Putting it differently, the “fingerprint” of a quantum phase, a quantum phase transition, and the underlying model dictating these phases and transitions are unique. The functional forms of $\overline{\mathcal{F}}$ and $\overline{\mathcal{D}}$ as we drive the system around its phase space are specific for a given model. A systematic study $\overline{\mathcal{F}}$ and $\overline{\mathcal{D}}$ enables the proper detection of the QCP using finite $T$ data and, interestingly, also the discovery of the underlying spin chain model giving rise to those quantum phases and quantum phase transitions.
We should also remark that the above features are still present for spin chains away from the thermodynamic limit. Preliminary calculations using spin chains containing about 10 qubits already show that we can detect the QCPs with reasonable accuracy and using finite temperature data only pav24b . Moreover, in order to properly exclude finite size effects or spurious discontinuities in the derivatives of $\overline{\mathcal{F}}$ and $\overline{\mathcal{D}}$ that are not related to QCPs, both the external and internal teleportation based QCP detectors must be used pav24b . Only discontinuities in the derivatives that occur in both quantities at the same place are the ones marking a QCP. Spurious discontinuities or finite size effects not related to a QCP do not occur at the same spot in both quantities pav24b . In other words, the internal teleportation QCP detector not only behaves differently from the external one but it is crucial in the task of identifying finite size effects or spurious discontinuities in the fidelity and in the trace distance that are not associated with QCPs.